2. Show that p-1 E(+1) = 2a +(*)(4+1) k=0 l=0 Hint: Fix an 0 <I<p – 1, Ask youself how many k satisfy k2 = 1 mod p.
Answer #1

where k= X-O 00 X=0 criterion - Now Eulers kp us P-1 i=o axps ear where Lemmat : Let f (2) P = 40 + ax + ... + k & the degour Now So K=O =#1 are mod.p. Thus p-1 lt 2 l=0 But P-3 4 ans.coer on the un cas é is a=1 b=o c=1 (allett) 1) xx-las pfy sinc

https://math.stackexchange.com/questions/1220507/how-many-values-of-k-satisfy-left-frackp-right-left-frack1p/1221135#1221135